OpenStudy (anonymous):

Alfonso is 8 years older than Ashley. Five years ago Alfonso was three times as old as Ashley. How old is Ashley now? A. 9 B. 17 C. 8

OpenStudy (australopithecus):

This is a system of equations question. set two variables, x = Alfonso's age y = Ashley's age Write two equations given by the question: Alfonso is 8 years older than Ashley. <- this is your first equation Five years ago Alfonso was three times as old as Ashley <- this is your second equation Sub on equation into the other. For example if you had two equations: y + 3 = x x + 5 = y Sub one into the other and solve for the remaining variable: so, y + 3 = x becomes, (x + 5) + 3 = x solve for x, Sub x into one of the equations and solve for y. You can easily check your answer by subbing both variables in and solving both sides of one of the equations you should get two numbers that equal eachother

OpenStudy (anonymous):

so the answer is 9?